Mass of lead (II) oxide needed to deposit 20.561g of lead metal
1
Expert's answer
2020-11-26T05:22:48-0500

Pb+O2+PbO2

n=m/M

n (PbO2) = n (Pb)

M (Pb) = 207.2 g/mol

M (PbO2) = 239.2 g/mol

n (Pb) = n (PbO2) = 20.561/207.2 = 0.1 mol

m (PbO2) = 0.1 x 239.2 = 24 g